# ConsistI2V: Enhancing Visual Consistency for Image-to-Video Generation

### v1

@@ Abstract
Image-to-video (I2V) generation aims to use the initial frame (alongside a text prompt) to create a video sequence. A grand challenge in I2V generation is to maintain visual consistency throughout the video: existing methods often struggle to preserve the integrity of the subject, background, and style from the first frame, as well as ensure a fluid and logical progression within the video narrative (cf. Figure 1). To mitigate these issues, we propose ConsistI2V, a diffusion-based method to enhance visual consistency for I2V generation. Specifically, we introduce (1) spatiotemporal attention over the first frame to maintain spatial and motion consistency, (2) noise initialization from the low-frequency band of the first frame to enhance layout consistency. These two approaches enable ConsistI2V to generate highly consistent videos. We also extend the proposed approaches to show their potential to improve consistency in auto-regressive long video generation and camera motion control. To verify the effectiveness of our method, we propose I2V-Bench, a comprehensive evaluation benchmark for I2V generation. Our automatic and human evaluation results demonstrate the superiority of ConsistI2V over existing methods.
Machine Learning, ICML
[…]
1 Introduction
[…]
@@ 1 Introduction
Figure 1: Comparison of image-to-video generation results obtained from SEINE (Chen et al., 2023c) and our ConsistI2V. SEINE shows degenerated appearance and motion as the video progresses, while our result maintains visual consistency. We feed the same first frame to SEINE and ConsistI2V and show the generated video on the right.
@@ 2 Related Work
Video Editing As paired video data before and after editing is hard to obtain, several methods (Ceylan et al., 2023; Wu et al., 2023b; Geyer et al., 2023; Wu et al., 2023a; Zhang et al., 2023b) employ a pretrained T2I model for zero-shot/few-shot video editing. To ensure temporal coherency between individual edited frames, these methods apply cross-frame attention mechanisms in the T2I model. Specifically, Tune-A-Video (Wu et al., 2023b) and Pix2Video (Ceylan et al., 2023) modify the T2I model’s self-attention layers to enable each frame to attend to its immediate previous frame and the video’s first frame. TokenFlow (Geyer et al., 2023) and Fairy (Wu et al., 2023a) select a set of anchor frames such that all anchor frames can attend to each other during self-attention.
Image-to-Video Generation Steering the video’s content using only text descriptions can be challenging. Recently, a myriad of methods utilizing both first frame and text for I2V generation have emerged as a solution to achieve more controllable video generation. Among these methods, Emu-Video (Girdhar et al., 2023), SEINE (Chen et al., 2023c), AnimateAnything (Dai et al., 2023) and PixelDance (Zeng et al., 2023) propose simple modifications to the T2V U-Net by concatenating the latent features of the first frame with input noise to enable first-frame conditioning. I2VGen-XL (Zhang et al., 2023a), Dynamicrafter (Xing et al., 2023) and Moonshot (Zhang et al., 2024) add extra image cross-attention layers in the model to inject stronger conditional signals into the generation process. Our approach varies from previous studies in two critical respects: (1) our spatiotemporal feature conditioning methods effectively leverage the input first frame, resulting in better visual consistency in the generated videos and enable efficient training on public video-text datasets.(2) We develop noise initialization strategies in the I2V inference processes, while previous I2V generation works rarely focus on this aspect.
@@ 3.4 Window-based Temporal Feature Conditioning
To effectively leverage the first frame features, we also augment the temporal self-attention layers to include a local window of features from the first frame. Formally, given an input hidden state $\mathbf{z}\in\mathbb{R}^{N\times C\times H\times W}$, where $N$ is the number of frames and $C,H,W$ correspond to the channel, height and width dimension of the hidden state tensor, our model first reshape its height and width dimension to the batch dimension to form a new hidden state $\mathbf{\bar{z}}\in\mathbb{R}^{(H\times W)\times N\times C}$. We then compute the query, key and value from $\mathbf{\bar{z}}$ as:
@@
where $\tilde{z}^{1}\in\mathbb{R}^{(H\times W)\times(K\times K-1)\times C}$ is a tensor constructed in a way that its $h\times w$ position in the batch dimension corresponds to an $K\times K$ window of the first frame features, centred at the spatial position of $(h,w)$. The first frame feature vector at $(h,w)$ is not included in $\tilde{z}^{1}$ as it is already presented in $\mathbf{\bar{z}}$. We pad $\mathbf{z}$ in its spatial dimensions by replicating the boundary values to ensure that all spatial positions in $\mathbf{z}$ will have a complete window around them. We then concatenate $\mathbf{\bar{z}}$ with $\tilde{z}^{1}$ to enlarge the sequence length of the key and value matrices. See Figure 3 to visualize the augmented temporal self-attention. Our main rationale is that the visual objects in a video may move to different spatial locations as the video progresses. Therefore, having an extra window of keys and values around the query location increases the probability of attending to the same entity in the first frame when performing temporal self-attention. In practice, we set $K=3$ to control the time complexity of the attention operations.
@@ 3.5 Inference-time Layout-Guided Noise Initialization
Figure 4: Visualization of different frequency bands after spatiotemporal frequency decomposition from the original latent $\mathbf{z}_{0}$.
Existing literature (Lin et al., 2024) in image diffusion models has identified that there exists a noise initialization gap between training and inference, due to the fact that common diffusion noise schedules create an information leak to the diffusion noise during training, causing it to be inconsistent with the random Gaussian noise sampled during inference. In the domain of video generation, this initialization gap has been further explored by FreeInit (Wu et al., 2023c), showing that the information leak mainly comes from the low-frequency component of a video after spatiotemporal frequency decomposition, and adding this low-frequency component to the initial inference noise greatly enhances the quality of the generated videos.
To better understand how different frequency bands correspond to different visual features in the video, we visualize the videos decoded from the VAE latents after spatiotemporal frequency decomposition, as shown in Figure 4. We observe that the video’s high-frequency component captures the fast-moving objects and the fine details in the video, whereas the low-frequency component corresponds to those slowly moving parts and represents an overall layout in each frame. Based on this observation, we propose FrameInit, which duplicates the input first frame into a static video and uses its low-frequency component as a coarse layout guidance during inference. Formally, given the latent representation $\mathbf{z}_{0}$ of a static video and an inference noise $\mathbf{\epsilon}$, we first add $\tau$ step inference noise to the static video to obtain $\mathbf{z}_{\tau}=\texttt{add\_noise}(\mathbf{z}_{0},\mathbf{\epsilon},\tau)$. We then extract the low-frequency component of $\mathbf{z}_{\tau}$ and mix it with $\mathbf{\epsilon}$:
@@ 4 I2V-Bench
Following VBench (Huang et al., 2023b), our evaluation framework encompasses two key dimensions, each addressing distinct aspects of Image-to-Video (I2V) performance: (1) Visual Quality assesses the perceptual quality of the video output regardless of the input prompts. We measure the subject and background consistency, temporal flickering, motion smoothness and dynamic degree. (2) Visual Consistency evaluates the video’s adherence to the text prompt given by the user. We measure object consistency, scene consistency and overall video-text consistency. Further details can be found in the Appendix C.
@@
UCF-101 MSR-VTT Human Evaluation - Consistency
@@
Table 2: Automatic evaluation results on I2V-Bench. Consist. denotes the consistency metrics. Bold: best results (excluding AnimateAnything). Underline: top-2 (excluding AnimateAnything).
@@
SEINE 96.08 97.92 42.12 93.98 88.21 35.43 27.68 20.21
ConsistI2V 96.65 97.77 37.48 94.69 90.85 32.06 24.81 19.50
5 Experiments
5.1 Implementation Details
5.2 Quantitative Evaluation
UCF-101 & MSR-VTT
@@ 5.3 Qualitative Evaluation
We present a visual comparison of our model with DynamiCrafter, SEINE and AnimateAnything in Figure 5. We exclude I2VGen-XL in this section as its generated video cannot fully adhere to the visual details from the input first frame. As shown in the figure, current methods often struggle with maintaining appearance and motion consistency in video sequences. This can include (1) sudden changes in subject appearance mid-video, as demonstrated in the “ice cream” case for DynamiCrafter and SEINE; (2) Background inconsistency, as observed in DynamiCrafter’s “biking” case; (3) unnatural object movements, evident in the “dog swimming” case (DynamiCrafter) and “tornado in a jar” case (SEINE) and (4) minimal or absent movement, as displayed in most of the generated videos by AnimateAnything. On the other hand, ConsistI2V produces videos with subjects that consistently align with the input first frame. Additionally, our generated videos exhibit more natural and logical motion, avoiding abrupt changes and thereby ensuring improved appearance and motion consistency. More visual results of our model can be found in Appendix H.
@@ 5.4 Ablation Studies
Table 3: Ablation study results for our proposed spatiotemporal first frame conditioning mechanisms and FrameInit on UCF-101. T.Cond. and S.Cond. correspond to temporal and spatial first frame conditioning, respectively.
FVD $\downarrow$ IS $\uparrow$ FID $\downarrow$
ConsistI2V 177.66 56.22 15.74
w/o FrameInit 245.79 42.21 24.08
w/o FrameInit & T.Cond. 224.16 42.98 24.04
w/o FrameInit & T.Cond & S.Cond. 704.48 21.97 68.39
@@
Spatiotemporal First Frame Conditioning Our ablation results on UCF-101 (c.f. Table 3) reflects a significant performance boost after applying the proposed spatial and temporal first frame conditioning mechanisms to our model. Although removing temporal first frame conditioning leads to an overall better performance for the three quantitative metrics, in practice we find that only using spatial conditioning often results in jittering motion and larger object distortions, evident in the last two frames in the third row of Figure 6. When both spatial and temporal first frame conditioning is removed, our model loses the capability of maintaining the appearance of the input first frame.
@@ 5.5 More Applications
While our image animation model provides native support for long video generation by reusing the last frame of the previous video to generate the subsequent video, we observe that directly using the model to generate long videos may lead to suboptimal results, as the artifacts in the previous video clip will often accumulate throughout the autoregressive generation process. We find that using FrameInit to guide the generation of each video chunk helps stabilize the autoregressive video generation process and results in a more consistent visual appearance throughout the video, as shown in Figure 7.
@@
When adopting FrameInit for inference, instead of using the static first frame video as the input, we can alternatively create synthetic camera motions from the first frame and use it as the layout condition. For instance, camera panning can be simulated by creating spatial crops in the first frame starting from one side and gradually moving to the other side. As shown in Figure 7, by simply tweaking the FrameInit parameters to $\tau=750$ and $D_{0}=0.5$ and using the synthetic camera motion as the layout guidance, we are able to achieve camera panning and zoom-in/zoom-out effects in the generated videos.
@@ A.2 Training Paradigms
Existing I2V generation models (Xing et al., 2023; Zeng et al., 2023; Zhang et al., 2023a, 2024) often employ joint video-image training (Ho et al., 2022b) that trains the model on video-text and image-text data in an interleaving fashion, or apply multi-stage training strategies that iteratively pretrain different model components using different types of data. Our model introduces two benefits over prior methods: (1) the explicit conditioning mechanism in the spatial and temporal self-attention layers effectively utilizes the visual cues from the first frame to render subsequent frames, thus reducing the difficulty of generating high-quality frames for the video diffusion model. (2) We directly employ the LDM VAE features as the conditional signal, avoiding training additional adaptor layers for other feature modalities (e.g. CLIP (Radford et al., 2021) image embeddings), which are often trained in separate stages by other methods. As a result, we train our model with a single video-text dataset in one stage, where we finetune all the parameters during training.

### v2

@@ Abstract
Image-to-video (I2V) generation aims to use the initial frame (alongside a text prompt) to create a video sequence. A grand challenge in I2V generation is to maintain visual consistency throughout the video: existing methods often struggle to preserve the integrity of the subject, background, and style from the first frame, as well as ensure a fluid and logical progression within the video narrative (cf. Figure 1). To mitigate these issues, we propose ConsistI2V  [^1] Project Website: https://tiger-ai-lab.github.io/ConsistI2V/, a diffusion-based method to enhance visual consistency for I2V generation. Specifically, we introduce (1) spatiotemporal attention over the first frame to maintain spatial and motion consistency, (2) noise initialization from the low-frequency band of the first frame to enhance layout consistency. These two approaches enable ConsistI2V to generate highly consistent videos. We also extend the proposed approaches to show their potential to improve consistency in auto-regressive long video generation and camera motion control. To verify the effectiveness of our method, we propose I2V-Bench, a comprehensive evaluation benchmark for I2V generation. Our automatic and human evaluation results demonstrate the superiority of ConsistI2V over existing methods.
@@
Figure 1: Comparison of image-to-video generation results obtained from SEINE (Chen et al., 2023c) and our ConsistI2V. SEINE shows degenerated appearance and motion as the video progresses, while our result maintains visual consistency. We feed the same first frame to SEINE and ConsistI2V and show the generated videos on the right.
1 Introduction
@@ 2 Related Work
Video Editing As paired video data before and after editing is hard to obtain, several methods (Ceylan et al., 2023; Wu et al., 2023b; Geyer et al., 2023; Wu et al., 2023a; Zhang et al., 2023b; Cong et al., 2023) employ a pretrained T2I model for zero-shot/few-shot video editing. To ensure temporal coherency between individual edited frames, these methods apply cross-frame attention mechanisms in the T2I model. Specifically, Tune-A-Video (Wu et al., 2023b) and Pix2Video (Ceylan et al., 2023) modify the T2I model’s self-attention layers to enable each frame to attend to its immediate previous frame and the video’s first frame. TokenFlow (Geyer et al., 2023) and Fairy (Wu et al., 2023a) select a set of anchor frames such that all anchor frames can attend to each other during self-attention.
Image-to-Video Generation Steering the video’s content using only text descriptions can be challenging. Recently, a myriad of methods utilizing both first frame and text for I2V generation have emerged as a solution to achieve more controllable video generation. Among these methods, Emu-Video (Girdhar et al., 2023), SEINE (Chen et al., 2023c), AnimateAnything (Dai et al., 2023) and PixelDance (Zeng et al., 2023) propose simple modifications to the T2V U-Net by concatenating the latent features of the first frame with input noise to enable first-frame conditioning. I2VGen-XL (Zhang et al., 2023a), Dynamicrafter (Xing et al., 2023) and Moonshot (Zhang et al., 2024) add extra image cross-attention layers in the model to inject stronger conditional signals into the generation process. Our approach varies from previous studies in two critical respects: (1) our spatiotemporal feature conditioning methods effectively leverage the input first frame, resulting in better visual consistency in the generated videos and enabling efficient training on public video-text datasets. (2) We develop noise initialization strategies in the I2V inference processes, while previous I2V generation works rarely focus on this aspect.
@@ 3.4 Window-based Temporal Feature Conditioning
[…]
To effectively leverage the first frame features, we also augment the temporal self-attention layers to include a broader window of features from the first frame. We propose to compute the query, key and value from $\mathbf{\bar{z}}$ as:
@@
where $\tilde{z}^{1}\in\mathbb{R}^{(H\times W)\times(K\times K-1)\times C}$ is a tensor constructed in a way that its $h\times w$ position in the batch dimension corresponds to an $K\times K$ window of the first frame features, centred at the spatial position of $(h,w)$. The first frame feature vector at $(h,w)$ is not included in $\tilde{z}^{1}$ as it is already presented in $\mathbf{\bar{z}}$, making the second dimension of $\tilde{z}^{1}$ as $K\times K-1$. We pad $\mathbf{z}$ in its spatial dimensions by replicating the boundary values to ensure that all spatial positions in $\mathbf{z}$ will have a complete window around them. We then concatenate $\mathbf{\bar{z}}$ with $\tilde{z}^{1}$ to enlarge the sequence length of the key and value matrices. See Figure 3 for a visualization of the augmented temporal self-attention. Our main rationale is that the visual objects in a video may move to different spatial locations as the video progresses. The vanilla or common formulation of temporal self-attention in video diffusion models essentially assumes $K=1$, as shown by the blue tokens in Figure 3. Having an extra window of keys and values around the query location with the window size $K>1$ (green tokens in Figure 3) increases the probability of attending to the same entity in the first frame when performing temporal self-attention. In practice, we set $K=3$ to control the time complexity of the attention operations. As the input latent is downsampled 8$\times$ by the VAE, and the denoising U-Net further downsamples the feature map in its intermediate layers, our selected window size ($K=3$) can still create a large receptive field in deeper U-Net layers. For example, in the second last level of the U-Net (32$\times$ total downsampling), a $256\times 256$ video will be compressed into a $8\times 8$ feature map. Consequently, a $3\times 3$ window on this feature map covers a region of $96\times 96$ in the input first frame. This is a significant increase from previous methods where a $1\times 1$ window on the feature map corresponded to only a $24\times 24$ region in the original frame.
@@ 3.5 Inference-time Layout-Guided Noise Initialization
Figure 4: Different frequency bands from the original latent $\mathbf{z}_{0}$ after spatiotemporal frequency decomposition.
Existing literature (Lin et al., 2024) in image diffusion models has identified that there exists a noise initialization gap between training and inference, due to the fact that common diffusion noise schedules create an information leak to the diffusion noise during training, causing it to be inconsistent with the random Gaussian noise sampled during inference. In the domain of video generation, this initialization gap has been further explored by FreeInit (Wu et al., 2023c), showing that the information leak mainly comes from the low-frequency component of a video after spatiotemporal frequency decomposition and adding this low-frequency component to the initial inference noise greatly enhances the quality of the generated videos. Inspired by the significance of frequency bands in images and videos for human perception, prior work like Blurring Diffusion Models (Hoogeboom & Salimans, 2022) employs diffusion processes in the frequency domain to enhance generation results. Here, we also aim to study how to effectively leverage the video’s different frequency components to enhance I2V generation.
To better understand how spatiotemporal frequency bands and visual features are correlated in videos, we visualize the videos decoded from the VAE latents after spatiotemporal frequency decomposition, as shown in Figure 4. We observe that the video’s high-frequency component captures the fast-moving objects and the fine details in the video, whereas the low-frequency component corresponds to those slowly moving parts and represents an overall layout. Based on this observation, we propose FrameInit, which duplicates the input first frame into a static video and uses its low-frequency component as a coarse layout guidance during inference. Formally, given the latent representation $\mathbf{z}_{0}$ of a static video and an inference noise $\mathbf{\epsilon}$, we first add $\tau$ step inference noise to the static video to obtain $\mathbf{z}_{\tau}=\texttt{add\_noise}(\mathbf{z}_{0},\mathbf{\epsilon},\tau)$. We then extract the low-frequency component of $\mathbf{z}_{\tau}$ and mix it with $\mathbf{\epsilon}$:
@@ 4 I2V-Bench
Following VBench (Huang et al., 2023b), our evaluation framework encompasses two key dimensions, each addressing distinct aspects of I2V performances: (1) Visual Quality assesses the perceptual quality of the video output regardless of the input prompts. We measure the subject and background consistency, temporal flickering, motion smoothness and dynamic degree. (2) Visual Consistency evaluates the video’s adherence to the text prompt given by the user. We measure object consistency, scene consistency and overall video-text consistency. Further details can be found in Appendix C.
5 Experiments
5.1 Implementation Details
5.2 Quantitative Evaluation
UCF-101 & MSR-VTT
@@ 5.2 Quantitative Evaluation
UCF-101 MSR-VTT Human Eval: Consistency
@@
Table 2: Automatic evaluation results on I2V-Bench. Consist. denotes the consistency metrics. Bold: best results. Underline: second best.
@@
[…]
@@ 5.3 Qualitative Evaluation
We present a visual comparison of our model with DynamiCrafter, SEINE and AnimateAnything in Figure 5. We exclude I2VGen-XL in this section as its generated video cannot fully adhere to the visual details from the input first frame. As shown in the figure, current methods often struggle with maintaining appearance and motion consistency in video sequences. This can include (1) sudden changes in subject appearance mid-video, as demonstrated in the “ice cream” case for DynamiCrafter and SEINE; (2) background inconsistency, as observed in DynamiCrafter’s “wooden figurine walking” case; (3) unnatural object movements, evident in the “dog swimming” case (DynamiCrafter) and “tornado in a jar” case (SEINE) and (4) minimal or absent movement, as displayed in most of the generated videos by AnimateAnything. On the other hand, ConsistI2V produces videos with subjects that consistently align with the input first frame. Additionally, our generated videos exhibit more natural and logical motion, avoiding abrupt changes and thereby ensuring improved appearance and motion consistency. More visual results of our model can be found in Appendix H.
@@ 5.4 Ablation Studies
Hyperparameters of FrameInit To further investigate how the FrameInit hyperparameters $\tau$ (number of steps of inference noise added to the first frame) and $D_{0}$ (spatiotemporal stop frequency) affect the generated videos, we conduct an additional qualitative experiment by generating the same video using different FrameInit hyperparameters and show the results in Figure 7. Comparing (1) and (4) in Figure 7, we find that while using our default setting of FrameInit achieves more stable outputs, it also slightly restricts the motion magnitude of the generated video compared to not using FrameInit at all. From comparisons between (1), (5), and (1), (6), we observe that either increasing $D_{0}$ alone or reducing $\tau$ alone does not significantly impact the motion magnitude of the videos. However, comparisons between (1), (2), and (3) reveal that when both $D_{0}$ is increased and $\tau$ is decreased, there is a further restriction in video dynamics. This suggests that the static first frame video exerts a stronger layout guidance, intensifying its influence on video dynamics.
@@
Table 3: UCF-101 ablation study results and runtime statistics for spatiotemporal first frame conditioning and FrameInit. T.Cond. and S.Cond. correspond to temporal and spatial first frame conditioning, respectively.
UCF-101 (Zero-Shot) Inference Cost Statistics
FVD $\downarrow$ IS $\uparrow$ FID $\downarrow$ GPU Mem. Runtime TFLOPs (UNet)
ConsistI2V 177.66 56.22 15.74 9249 Mb 18.48s 5.168
w/o FrameInit 245.79 42.21 24.08 9245 Mb 18.37s 5.168
w/o FrameInit & T.Cond. 224.16 42.98 24.04 9233 Mb 17.93s 5.117
w/o FrameInit & T.Cond. & S.Cond. 704.48 21.97 68.39 9017 Mb 17.14s 5.015
Figure 7: Visual comparisons of employing different FrameInit hyperparameters during I2V generation. All videos are generated using the same seed.
Spatiotemporal First Frame Conditioning Our ablation results on UCF-101 (c.f. Table 3) reflects a significant performance boost after applying the proposed spatial and temporal first frame conditioning mechanisms to our model. Although removing temporal first frame conditioning leads to an overall better performance for the three quantitative metrics, in practice we find that only using spatial conditioning often results in jittering motion and larger object distortions, evident in the third row of Figure 6. When both spatial and temporal first frame conditioning is removed, our model loses the capability of maintaining the appearance of the input first frame.
Runtime Efficiency To better understand how the proposed model components affect the inference cost, we also benchmark the runtime statistics of our ConsistI2V and its variants after disabling different components. We measure all models on a single Nvidia RTX 4090 with float32 and an inference batch size of 1. For TFLOPs computation, we only consider the denoising U-Net and exclude components such as the VAE and text encoder as these modules share the same computation across all model variants. As shown in the right part of Table 3, we find that the proposed spatiotemporal feature conditioning and FrameInit incur limited computational overhead during inference, introducing approximately 200Mb extra GPU memory and $\sim$1.3s additional inference time. These results demonstrate that our method is highly efficient during inference.
@@ 5.5 More Applications
While our I2V model provides native support for long video generation by reusing the last frame of the previous video to generate the next video, we observe that directly using the model to generate long videos may lead to suboptimal results, as the artifacts in the previous video clip will often accumulate throughout the autoregressive generation process. We find that using FrameInit to guide the generation of each video chunk helps stabilize the autoregressive video generation process and results in a more consistent visual appearance throughout the video, as shown in Figure 8.
@@
When adopting FrameInit for inference, instead of using the static first frame video as the input, we can alternatively create synthetic camera motions from the first frame and use it as the layout condition. For instance, camera panning can be simulated by creating spatial crops in the first frame starting from one side and gradually moving to the other side. As shown in the first two examples under camera motion control in Figure 8, by simply tweaking the FrameInit parameters to $\tau=750$ and $D_{0}=0.5$ and using the synthetic camera motion as the layout guidance, we are able to achieve camera panning and zoom-in/zoom-out effects in the generated videos without any additional training. For camera motions that involve perspective view change (e.g. rotating), we propose to apply image warping to the input first frame to create a sequence of perspective transformations. We then use this synthetic perspective view change as the layout condition to generate new videos. As shown in the last example of Figure 8, we can synthesize camera rotation by creating synthetic perspective view changes based on the coordinates of four selected key points in the first frame. As perspective view changes are often more complex than 2D panning/zooming, the FrameInit hyperparameters for this example are selected as $\tau=700$ and $D_{0}=0.5$ to enforce a stronger layout guidance.
6 Limitations
@@ A.2 Training Paradigms
Existing I2V generation models (Xing et al., 2023; Zeng et al., 2023; Zhang et al., 2023a; 2024) often employ joint video-image training (Ho et al., 2022b) that trains the model on video-text and image-text data in an interleaving fashion, or apply multi-stage training strategies that iteratively pretrain different model components using different types of data. Our model introduces two benefits over prior methods: (1) the explicit conditioning mechanism in the spatial and temporal self-attention layers effectively utilizes the visual cues from the first frame to render subsequent frames, thus reducing the difficulty of generating high-quality frames for the video diffusion model. (2) We directly employ the LDM VAE features as the conditional signal, avoiding training additional adaptor layers for other feature modalities (e.g. CLIP (Radford et al., 2021) image embeddings), which are often trained in separate stages by other methods. As a result, we train our model with a single video-text dataset in one stage, where we finetune all the parameters during training.